OpenStudy (anonymous):

\[\sum_{n=1}^{\infty}3(2/3)^(n+1)\]

OpenStudy (amistre64):

that was a good effort :)

OpenStudy (amistre64):

gonna have to prolly rewrite it tho

OpenStudy (anonymous):

is this \[\sum3^{\frac{2}{3}}(n+1)\]?

OpenStudy (anonymous):

no probably not because that would be infinite

OpenStudy (anonymous):

ahhhhk a geometric series! ignore the first 3 or rather pull it out front of the summation

OpenStudy (anonymous):

6

OpenStudy (anonymous):

just 6?

OpenStudy (anonymous):

yeah

OpenStudy (anonymous):

actually i think it is 4

OpenStudy (anonymous):

soooo 4?

OpenStudy (anonymous):

sorry 4...lol

OpenStudy (anonymous):

0 or 4?

OpenStudy (anonymous):

\[\sum_{n=1}^{\infty} (\frac{2}{3})^{n+1}\]=

OpenStudy (anonymous):

4

OpenStudy (anonymous):

change n + 1 to n and start at n = 0

OpenStudy (amistre64):

.... i change my vote to 'im an idiot' :)

OpenStudy (anonymous):

ok so 4 it is

OpenStudy (anonymous):

it is 4 if you just want the answer

OpenStudy (anonymous):

yesssss

OpenStudy (anonymous):

thnx

OpenStudy (anonymous):

NO THANKK YOUUUU :)

OpenStudy (anonymous):

actually satellite73, if you change n+1 to n you would have to start from n=2